OpenStudy (anonymous):

The potential energy of a charge q moving from A to B at point A is 5.4 × 10^-10 joules and at point B is 2.3 × 10^-10 joules. The charge is 2.5 × 10^-15 coulombs. What potential difference does the particle undergo? A. 1.2 × 10^3 volts B. 2.3 × 10^3 volts C. 2.5 × 10^4 volts D. 5.4 × 10^4 volts E. 7.0 × 10^4 volts

OpenStudy (anonymous):

\[V = \frac{U}{q}\] so if you want the change in potential difference.... you have \[\Delta V = \frac{\Delta U}{q} = \frac{U_{b} - U_{a}}{q} = \]

OpenStudy (anonymous):

You can do the rest!

OpenStudy (anonymous):

whats the U & Q stand for ?